Litho Morning Status 10/27

Gary Sosa gsosa at stanford.edu
Wed Oct 27 07:19:50 PDT 2010


Good Morning All....

   Here is the Litho morning status report.

Karl Suss 1: User accidentally depressed the EMO. Will power up 
tool/lamp and do a full hardware
reset and lamp check.

SVG Dev 1: Wafer transfer problem. Will investigate and repair as needed

EV Bonder: User reports top electrode temperature problem. Will 
investigate. May need to replace the
heater lamps.

   Please check back in Coral for current status and updates.

  Thanks....  The Litho Team



*/Photolithography Area Status/*

	

	

*/Wednesday, October 27, 2010/*

	

	

	

	

*/General Information / Lab Activities/*

	

	

	

Mario on vacation today

Staff IRT meeting- 9:00 AM

*/Exposure Tools/*

	

Status

	

Comments

	

Priority

ASML Stepper

	

*UP*

	

	

EVAlign

	

*UP*

	

	

EVAlign2

	

*UP*

	

	

EVG-Imprint

	

*UP*

	

	

Karl Suss 1

	

*DOWN*

	

EMO Accidentally depressed

	

1

Karl Suss 2

	

*PROBLEM*

	

	

*/Manual Coaters/*

	

Status

	

	

EVGSpraycoat

	

*PROBLEM*

	

Wobbly chuck

	

5

Headway

	

*UP*

	

	

Laurel

	

*UP*

	

	

*/Coaters/*

	

Process Status

	

	

SVGCoat

	

3612 Process

	

*Up*

	

	

3617 Process

	

*Up*

	

	

220-7 Process

	

*Problem*

	

Optimizing process

	

4

SVGCoat2

	

3612 Process

	

*Up*

	

	

955-0.7 Process

	

*Up*

	

	

220-3 Process

	

*Up*

	

	

*/Developers/*

	

Status

	

	

SVGDEV

	

*DOWN*

	

Wafer transfer problem

	

2

SVGDEV2

	

*UP*

	

	

*/Ovens / Hotplates/*

	

Status

	

	

BLUEM

	

*UP*

	

	

Blue M 90 C.

	

*UP*

	

	

Blue M 110 C

	

*PROBLEM*

	

Temperature Drift. Controller on order

	

7

Blue M 150 C

	

*PROBLEM*

	

Temperature Drift. Controller on order

	

8

Fusion

	

*DOWN*

	

Hotplate experiments

	

UV-Cure

	

*UP*

	

	

White-Oven

	

*UP*

	

	

YES Vapor Prime

	

*UP*

	

	

*/Hotplates- /*

	

	

	

Laurel-left

	

*UP*

	

	

Laurel- right

	

*UP*

	

	

Headway

	

*UP*

	

	

Torrey Pines- left

	

*UP*

	

	

90 Deg C

	

*UP*

	

	

115 Deg C

	

*UP*

	

	

Torrey Pines- Right

	

*UP*

	

	

*/Metrology/Inspection/*

	

Status

	

	

Alpha-Step 500

	

*UP*

	

	

Nanospec 2-Litho

	

*UP*

	

	

Nanospec 1-Etch

	

*UP*

	

	

P2 Profiler

	

*PROBLEM*

	

Error -11 on Unload, reset to recover

	

6

*/Microscpoes-/*

	

	

	

#1 Leica/Analytical

	

*UP*

	

	

#2 Ergolux/Drytek2

	

*UP*

	

	

#3 Olympus/Litho

	

*UP*

	

Camera App and Data OK

	

#4 Nikon Optiphot/Litho

	

*UP*

	

	

#5 Nikon/Litho

	

*UP*

	

	

#6 Leitz/Bosch

	

*UP*

	

	

#7 Leitz/Diffusion

	

*UP*

	

	

#8 Olympus/Litho

	

*UP*

	

	

#9 Olympus/Litho

	

*UP*

	

	

*/Miscellaneous/*

	

Status

	

	

EVBond

	

*PROBLEM*

	

Top Heater problem

	

3

KSBonder

	

*UP*

	

	

KSCoat

	

*UP*

	

	

Litho SRD

	

*UP*

	

	

	

	

	

	

	

	


-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/litho/attachments/20101027/6c4cb6e3/attachment.html>


More information about the litho mailing list